# Augmented Lagrangian Solvers for Poroelasticity with Fracture Contact Mechanics

### v1

@@ Abstract
In the subsurface, fractures and the surrounding porous rock can deform in interaction with fluid flow. Advanced mathematical models governing these coupled processes typically combine fluid flow, poroelasticity, and fracture contact mechanics, representing fractures as co-dimension one objects within the porous medium. The resulting system of equations is complex and highly nonlinear. As a result, convergence issues with nonlinear solvers are common, which hinders the practical implementation of such models.
One particular source of difficulty for the nonlinear solvers comes from the fracture contact mechanics, due to its inherently nonsmooth character. In this paper, we investigate solvers based on the augmented Lagrangian formulation of the frictional contact problem. This includes two classical solvers, namely the generalized Newton method (using complementarity functions) and the return map method. In addition, we propose a new augmented Lagrangian solver that combines the two approaches.
Numerical experiments in two and three dimensions are conducted to assess the performance of the solvers on problems of mixed-dimensional poromechanics with fracture contact mechanics. The experiments indicate that for simpler setups, the solvers behave quite predictably and in accordance with established heuristics from the contact mechanics literature. However, as the complexity of the problem increases, the solvers become more unpredictable, and break with these heuristics. In particular, the two classical solvers become highly sensitive to the augmentation parameter. The new solver converges across a larger range of this parameter in most cases, but nevertheless also displays unpredictable behavior.
@@
Fractured porous media, poroelasticity, contact mechanics, augmented Lagrangian, generalized Newton, return map
@@ 1 Introduction
An overview of different conceptual approaches to model fractured porous media can be found in [1, 2, 3]. In this work, we specifically consider discrete fracture-matrix models, in which the fractures are explicitly represented. Specifically, the matrix (i.e. the porous medium surrounding the fractures) is modeled as linearly poroelastic [4] with flow governed by Darcy’s law. For the fracture, Darcy flow is also considered, while deformation is modeled by contact mechanics including Coulomb friction [5]. There have been several recent works on such models [6, 7, 8, 9, 10, 11, 12, 13, 14], focusing on the development of discretization methods adapted to the discrete fracture-matrix model. In this context, for the mechanical subproblem, a common approach is to use finite element methods [10, 11, 12, 13, 14], but, motivated by the coupling with flow, finite volume methods are also increasingly being applied [6, 7, 8, 9]. The flow equations are usually discretized by finite volume methods [6, 7, 8, 9, 10, 11, 12, 13], but finite element methods have also been used [14]. In addition, considerable efforts have been made to stabilize the mechanical subproblem [15, 12, 16].
@@
To apply these discretization schemes to complex setups that reflect the physical states and fractured structures encountered in field applications, there is a need for nonlinear solvers for discrete fracture-matrix models for poromechanics, and scalable linear solvers. Linear solvers, based on block preconditioning that is tailored to both the individual physical processes and their couplings, have been developed for both finite element [17, 18, 19] and finite volume [20, 21] discretizations. In terms of nonlinear solvers, a challenging aspect is the inherent nonlinear and nonsmooth character of frictional contact problems [5, 22]. Several classes of nonlinear solvers that were originally developed for frictional contact mechanics have been extended to account for poromechanics. These solvers are built around different ways of formulating the contact conditions, and includes solvers based on the penalty [10], augmented Lagrangian [6, 23] and Nitsche formulations [11, 12] of the frictional contact problem. However, the poromechanical coupling introduces new challenges for the nonlinear solvers: First, the extra couplings make for larger and more complex algebraic systems. Second, depending on model choices, the systems might contain other nonlinearities that are triggered by the contact mechanics, such as shear dilation of fractures and fracture permeability modeled by a cubic law [7, 8]. As a result, the design of robust nonlinear solvers for these problems remains a challenge that must be addressed to enable reliable simulations with advanced physical models on complex fracture geometries.
The focus of this paper is on nonlinear solvers based on the augmented Lagrangian formulation of frictional contact mechanics. It may be viewed as a combination of the Lagrange multiplier and penalty formulations, inheriting favorable properties from both. An important feature of the augmented Lagrangian formulation, as well as the penalty formulation, is the presence of a numerical parameter which must be chosen by the user (we will henceforth refer to it as the augmentation parameter). This parameter can be interpreted as regularizing the problem, although the solution to the augmented Lagrangian formulation will still be exact for any positive value of the parameter. This is in contrast to the penalty formulation, where the original solution is only recovered as the augmentation parameter goes to infinity. While the augmented Lagrangian formulation is a popular choice for solving problems involving contact mechanics, it is still prone to convergence issues. This is demonstrated in [24], where several nonlinear solvers, among which solvers based on the augmented Lagrangian formulation, were tested on a range of purely mechanical contact problems for non-fractured media. The solvers failed to converge in several cases, with the performance often being very sensitive to the value of the augmentation parameter. The present work can be viewed as an extension of the parts of the study in [24] concerning the augmented Lagrangian solvers, to assess the performance of these solvers on problems of coupled flow and deformation including fracture contact mechanics.
@@
There are two distinct classical nonlinear solvers based on the augmented Lagrangian formulation of frictional contact mechanics. The first approach is based on a reformulation of the variational inequalities of contact mechanics as the zero sets of certain non-smooth functions called complementarity functions. This enables the use of a generalized Newton method with appropriately defined generalized Jacobians to handle the system’s non-smoothness. This approach was first introduced by Alart and Curnier [25], and several different ways of formulating the complementarity functions have since been derived [26, 27]. The second approach, introduced by Simo and Laursen [28], uses a fixed-point iteration on the augmented Lagrange multipliers, resulting in a return map method that can also be interpreted as an Uzawa algorithm [24]. We note that the algorithm in [28] corresponds to an implicit version of the Uzawa algorithm, which is known to improve the stability of the algorithm [29], at the cost of needing to solve a more complicated system.
@@
Both the generalized Newton method and the return map method will be considered in this work. Moreover, we will provide a formal connection between these two methods. More specifically, we will demonstrate that the return map method is equivalent to solving a sequence of nonlinear systems with regularized complementarity functions, where each of these systems can be solved by the generalized Newton method. This observation leads us to propose a new nonlinear solver, in which we incorporate the return map into the generalized Newton method as a postprocessing step. This approach allows the return map to be applied without requiring regularization of the equations.
@@
To illustrate the convergence challenges and evaluate the performance of the generalized Newton method, the return map method, and our new nonlinear solver, we carry out a series of numerical experiments in both two and three dimensions. These experiments are inspired by hydraulic stimulation of geothermal reservoirs, in which critically stressed fractures are stimulated by fluid injection, resulting in fracture slip and opening. Simulations are conducted across a range of injection pressures, as well as a range of shear dilation angles, highlighting how the coupling between flow and mechanics evolves and contributes to convergence difficulties for the nonlinear solvers. We also investigate a range of values for the augmentation parameter to assess whether established heuristics from contact mechanics can inform its selection in this context.
@@
The remainder of the paper is structured as follows. In Section 2, we present the mathematical model for coupled flow and deformation in fractured porous media. Section 3 introduces the generalized Newton method and the return map method, adapted to the governing equations, and establishes a formal connection between them. This connection motivates the development of our new nonlinear solver, which is subsequently introduced. In Section 4, we evaluate the performance of all three nonlinear solvers through a series of numerical experiments. Concluding remarks are provided in Section 5.
@@ 2 Mathematical model
We consider a discrete fracture-matrix model, where the fractures are modeled as co-dimension one objects, leading to a mixed-dimensional poromechanical model including fracture contact mechanics. The presentation of the model in this section closely follows that of [8]; see also [6, 30, 7]. We first give an overview of the mixed-dimensional geometry in section 2.1. In section 2.2, conservation laws for mass and momentum balance, with corresponding constitutive laws, are presented. The equations of fracture contact mechanics are presented in section 2.3. Finally, in section 2.4 we define the boundary conditions and provide a summary of the primary variables and equations.
@@ 2.1 Mixed-dimensional geometry
The fractured porous medium is described as a collection of subdomains $\Omega_{i}$ of varying dimension $d_{i}\in\{0,...,D\}$, with $D\in\{2,3\}$. The matrix subdomain will have dimension $D$, while fractures are modeled as co-dimension one objects, and will accordingly have dimension D1D-1. Moreover, intersections of fractures will have dimension D2D-2, and in the case of $D=3$, one also considers intersections of fracture intersection lines as zero-dimensional subdomains. The width of a fracture is characterized by its aperture $a_{i}$. This is further generalized by introducing the specific volume $\mathcal{V}_{i}=a_{i}^{D-d_{i}}$, accounting for dimension reduction for a subdomain $\Omega_{i}$ with any dimension $d_{i}\leq D$.
@@
Figure 1: Illustration of the mixed-dimensional geometry. The higher-dimensional subdomain $\Omega_{h}$ is coupled to the lower-dimensional subdomain $\Omega_{l}$ through the interfaces $\Gamma_{j}$ and $\Gamma_{k}$. Note that $\partial_{j}\Omega_{h},\partial_{k}\Omega_{h},\Gamma_{j},\Gamma_{k}$ and $\Omega_{l}$ all coincide geometrically. Figure adapted from [7]
@@ 2.2 Conservation laws and constitutive relations
Next, we define constitutive relations for the various quantities defined above, most of which are derived from [4]. The volumetric fluid flux is modeled by Darcy’s law:
@@
while in fracture subdomains, it is given by a cubic law [31]:
$\bm{\mathcal{K}}_{i}=\frac{a_{i}^{2}}{12}\mathbf{I}_{D-1}\ \ ,\ \ d_{i}=D-1.$ (5)
Finally, intersection permeability is computed as the average permeability of the intersecting fractures. Solid density is assumed to be constant, while the density of the slightly compressible fluid is given by (we omit the subscript $i$ for readability):
@@
where $\gamma$ is the compressibility of the fluid. The subscript ”ref” is used to denote reference values of variables. The interface flux $v_{j}$ is given by an averaged Darcy’s law [32]:
@@
Here, $\bm{n}_{h}$ is the outward normal vector of the higher-dimensional neighboring subdomain. The interface permeability $\mathcal{K}_{j}$ is also modeled by a cubic law, inheriting the aperture from its lower-dimensional neighboring subdomain,
$\mathcal{K}_{j}=\frac{\Pi_{j}^{l}a_{l}^{2}}{12},$ (8)
@@
where $\rho^{s}$ is the solid density. The only constitutive law now remaining is the one for the aperture $a_{i}$; this will be defined in section 2.3, after some new notation has been introduced.
@@ 2.3 Fracture contact mechanics
The equations of contact mechanics are imposed on pairs of $D$-dimensional matrix subdomains $\Omega_{h}$ and D1D-1-dimensional fracture subdomains $\Omega_{l}$ that are connected through two interfaces $\Gamma_{j},\Gamma_{k}$, as illustrated in figure 1. Some new variables are introduced: The fracture contact traction $\bm{\lambda}_{l}$, defined on the fracture domain $\Omega_{l}$, and the displacements $\bm{u}_{j},\bm{u}_{k}$ of the interfaces $\Gamma_{j},\Gamma_{k}$. For readability purposes, we will omit the subscript $l$ from the contact traction. The forces on the two fracture surfaces must be balanced according to Newton’s third law. Taking into account that the pressure from the fluid in the fracture contributes to the total fracture traction, the following balance of forces are imposed at the interfaces:
@@
$\begin{cases}\text{If}\ \lambda_{n}=0\implies\bm{\lambda}_{\tau}=0.\\ \text{Else}:\\ \ \ \begin{aligned} \lVert\bm{\lambda}_{\tau}\rVert&\leq-F\lambda_{n},\\ \lVert\bm{\lambda}_{\tau}\rVert&<-F\lambda_{n}\implies[[\dot{\bm{u}}]]_{\tau}=\bm{0},\\ \lVert\bm{\lambda}_{\tau}\rVert&=-F\lambda_{n}\implies\exists\ \zeta\in\mathbb{R}^{+}:[[\dot{\bm{u}}]]_{\tau}=\zeta\bm{\lambda}_{\tau},\end{aligned}\end{cases}$ (17)
where $F$ is the coefficient of friction and $\dot{\bm{u}}]]_{\tau}$ is the time derivative of the tangential displacement jump. We will for convenience refer to the entirety of equation (17) as the Coulomb friction law. Finally, we define the constitutive law for the aperture. In fracture subdomains, we set the aperture to equal the normal component of the interface displacement jump, with an additional constant reference aperture $a_{\mathrm{ref}}$ taking into account small-scale roughnesses in the undeformed state:
$a=a_{\mathrm{ref}}+[[\bm{u}]]_{n}.$ (18)
For intersection subdomains $(d_{i}<D-1)$, the aperture is computed as the average aperture of the intersecting fractures:
$a_{i}=\frac{1}{\lvert\hat{S}_{i}\rvert}\sum_{j\in\hat{S}_{i}}\Xi_{j}^{i}\Pi_{j}^{h}a_{h}.$ (19)

### v2

# arXiv:2508.11508v2[math.NA] 01 Mar 2026

@@ Abstract
In the subsurface, fractures and the surrounding porous rock can deform in interaction with fluid flow. Advanced mathematical models governing these coupled processes typically combine fluid flow, poroelasticity, and fracture contact mechanics. The resulting system of equations is complex and highly nonlinear. As a result, convergence issues with nonlinear solvers are common, causing a bottleneck for the numerical solution of such models.
One source of difficulty for the nonlinear solvers comes from the fracture contact mechanics, due to its inherently nonsmooth character. In addition, depending on the chosen constitutive model, the degree of nonlinearity is increased through coupling of flow and contact mechanics. In this paper, we investigate solvers based on the augmented Lagrangian formulation of the frictional contact problem. This includes two classical solvers, namely the generalized Newton method (using complementarity functions) and the return map method (equivalent to an Uzawa method). In addition, we propose a new solver that combines features of both approaches.
Numerical experiments in two and three dimensions, designed to simulate hydraulic stimulation of geothermal reservoirs, are conducted to assess the performance of the solvers on problems of poromechanics with fracture contact mechanics. The return map method has more difficulty handling the nonlinear coupling between flow and contact mechanics than the other solvers, in many cases not converging or using an excessive number of iterations. Our new combined solver performs the most robustly across the experiments, its performance being less sensitive to the value of the augmentation parameter than the other solvers.
@@
Fractured porous media, poroelasticity, contact mechanics, augmented Lagrangian, generalized Newton, return map, Uzawa, projected Newton
@@ 1 Introduction
An overview of different conceptual approaches to model fractured porous media can be found in [fracture_review, viswanathan2022fluid, vaezi2025review]. In this work, we specifically consider discrete fracture-matrix models, in which the fractures are explicitly represented. Specifically, the matrix (i.e. the porous medium surrounding the fractures) is modeled as linearly poroelastic [coussy] with flow governed by Darcy’s law. For the fracture, Darcy flow is also considered, while deformation is modeled by contact mechanics including Coulomb friction [kikuchi_oden]. There have been several recent works on such models [berge, Stefansson_2021, stefansson2023flexible, NOVIKOV2022111598, garipov2016, GARIPOV2019104075, BEAUDE2023116124, droniou_masson, SALIMZADEH2018212], focusing on the development of discretization methods adapted to the discrete fracture-matrix model. In this context, for the mechanical subproblem, a common approach is to use finite element methods [garipov2016, GARIPOV2019104075, BEAUDE2023116124, droniou_masson, SALIMZADEH2018212], but, motivated by the coupling with flow, finite volume methods are also increasingly being applied [berge, Stefansson_2021, stefansson2023flexible, NOVIKOV2022111598]. The flow equations are commonly discretized by finite volume methods [berge, Stefansson_2021, stefansson2023flexible, NOVIKOV2022111598, garipov2016, GARIPOV2019104075, BEAUDE2023116124, droniou_masson], but finite element methods have also been used [SALIMZADEH2018212]. In addition, considerable efforts have been made to stabilize the mechanical subproblem [FRANCESCHINI2020113161, BEAUDE2023116124, uzawa_bubble].
@@
To apply these discretization schemes to complex setups that reflect the physical states and fractured structures encountered in field applications, there is a need for nonlinear solvers for discrete fracture-matrix models for poromechanics, and scalable linear solvers. Linear solvers, based on block preconditioning that is tailored to both the individual physical processes and their couplings, have been developed for both finite element [franceschini2019block, franceschini2022scalable, franceschini2022scalable2] and finite volume [zabegaev2025efficient, zabegaev2025block] discretizations. In terms of nonlinear solvers, a challenging aspect is the inherent nonlinear and nonsmooth character of frictional contact problems [kikuchi_oden, WriggersPeter2006CCM]. Several classes of nonlinear solvers that were originally developed for frictional contact mechanics have been extended to account for poromechanics. These solvers are built around different ways of formulating the contact conditions, and includes solvers based on the penalty [NOVIKOV2022111598, garipov2016], augmented Lagrangian [berge, Stefansson_2021, BEAUDE2023116124, SALIMZADEH2018212] and Nitsche formulations [GARIPOV2019104075, BEAUDE2023116124] of the frictional contact problem. However, the poromechanical coupling introduces new challenges for the nonlinear solvers. First, the extra couplings make for larger and more complex algebraic systems. Second, depending on model choices, the systems might contain other nonlinearities that are triggered by the contact mechanics, such as shear dilation of fractures and fracture permeability modeled by a cubic law [Stefansson_2021, stefansson2023flexible]. As a result, the design of robust nonlinear solvers for these problems remains a challenge that must be addressed to enable reliable simulations with advanced physical models on complex fracture geometries.
@@
Nonlinear solvers based on the augmented Lagrangian formulation of frictional contact mechanics typically follows one of two distinctly different approaches. The first approach is based on a reformulation of the variational inequalities of contact mechanics as the zero sets of certain non-smooth functions called complementarity functions. This enables the use of a generalized Newton method with appropriately defined generalized Jacobians to handle the system’s non-smoothness. This approach was first introduced by Alart and Curnier [ALART1991353], and several different ways of formulating the complementarity functions have since been derived [klarbring, hueber]. The second approach, introduced by Simo and Laursen [SIMO199297], uses a fixed-point iteration on the augmented Lagrange multipliers, resulting in an implicit return map method that can also be interpreted as an implicit Uzawa algorithm [RenardSurvey]. Both approaches are well-established in the contact mechanics literature, and a systematic comparison of these two solvers (in addition to other nonlinear solvers) has been conducted in [RenardSurvey] for problems of a linearly elastic body in contact with a rigid obstacle, but not considering fractured media.
The generalized Newton method based on complementarity functions has been used in several of the aforementioned works to solve problems of coupled flow and deformation including fracture contact mechanics [berge, Stefansson_2021, BEAUDE2023116124]. The implicit return map method, using the augmented Lagrangian formulation, appears to have been less used to solve such problems. It was used together with a pure penalty formulation in [garipov2016, NOVIKOV2022111598]. In [nejati2016finite], an augmented Lagrangian method was developed for purely mechanical contact problems in fractured media, where a return map was used to augment the fracture gap, rather than the Lagrange multipliers. The result is an algorithm that is distinctly different from the one by Simo and Laursen [SIMO199297]. It was subsequently used in [SALIMZADEH2018212] to solve a coupled thermo-hydro-mechanical model.
In the present work, we perform a systematic study and comparison of three different nonlinear solvers based on the augmented Lagrangian formulation. These include the generalized Newton method of Alart and Curnier [ALART1991353], using the complementarity functions of Hüeber et al. [hueber], and the implicit return map method of Simo and Laursen [SIMO199297]. A systematic comparison of these two solvers, in the vein of [RenardSurvey], is lacking for problems of coupled flow and deformation including fracture contact mechanics. The third solver in our study is new, and it combines features of both the previous two solvers. Specifically, it is a quasi-Newton method that incorporates the return map as a postprocessing step after every generalized Newton iteration. Conceptually, this solver can be interpreted as a type of projected Newton method, which is a method used to solve constrained optimization problems [projected_newton], where each Newton iteration is followed by a return map (projection) onto the feasible set. A comparable approach has been employed by Jha et al. [jha2014coupled] to address purely mechanical subproblems within a fixed-stress solver for coupled multiphase poromechanics in fractured media. However, to the authors’ knowledge, the present work is the first time this type of solver has been applied to solve coupled flow and mechanics in a monolithic manner.
@@
Moreover, we show a mathematical connection between the generalized Newton and the implicit return map methods. Specifically, we will demonstrate that the implicit return map method is equivalent to solving a sequence of nonlinear systems with regularized complementarity functions, where each of these systems can be solved by the generalized Newton method. To the authors’ knowledge, the literature lacks a detailed derivation of this connection, which is key to understanding how our new, combined solver differs from both the generalized Newton and the implicit return map methods.
@@
To illustrate the convergence challenges and evaluate the performance of the generalized Newton method, the implicit return map method, and our new nonlinear solver, we carry out a series of numerical experiments in both two and three dimensions. These experiments are inspired by hydraulic stimulation of geothermal reservoirs, in which critically stressed fractures are stimulated by fluid injection, resulting in fracture slip and corresponding shear dilation. In the simulations, we use different constitutive model choices for the relationship between slip, aperture dilation and fracture permeability, to gradually increase the degree of nonlinear couplings between flow and mechanics, highlighting how such couplings can contribute to convergence difficulties. We also investigate a range of values for the so-called augmentation parameter; this is a numerical parameter inherent to the augmented Lagrangian formulation, which must be chosen by the user. This is relevant to study, as the value of this parameter can affect the convergence behavior of the nonlinear solvers, as shown in the study done in [RenardSurvey] on purely mechanical contact problems.
@@
The remainder of the paper is structured as follows. In Section 2, we present the mathematical model for coupled flow and deformation in fractured porous media. Section 3 introduces the generalized Newton method and the implicit return map method, adapted to the governing equations, and establishes a formal connection between them. Our new nonlinear solver, which combines features of both previous solvers, is subsequently introduced. In Section 4, we evaluate the performance of all three nonlinear solvers through a series of numerical experiments. Concluding remarks are provided in Section 5.
@@ 2 Mathematical model
We consider a discrete fracture-matrix model, where the fractures are modeled as co-dimension one objects, leading to a mixed-dimensional poromechanical model including fracture contact mechanics. The presentation of the model in this section closely follows that of [stefansson2023flexible]; see also [berge, porepy, Stefansson_2021]. We first give an overview of the mixed-dimensional geometry in section 2.1. In section 2.2, conservation laws for mass and momentum balance, with corresponding constitutive laws, are presented. The equations of fracture contact mechanics are presented in section 2.3. Finally, in section 2.4 we define the boundary conditions and provide a summary of the primary variables and equations.
@@ 2.1 Mixed-dimensional geometry
The fractured porous medium is described as a collection of subdomains $\Omega_{i}$ of varying dimension $d_{i}\in\{0,...,D\}$, with $D\in\{2,3\}$. The matrix subdomain will have dimension $D$, while fractures are modeled as co-dimension one objects, and will accordingly have dimension D1D-1. Moreover, intersections of fractures will have dimension D2D-2, and in the case of $D=3$, one also considers intersections of fracture intersection lines as zero-dimensional subdomains. The width of a fracture is characterized by its volumetric aperture $a_{i}$. This is further generalized by introducing the specific volume,
$\mathcal{V}_{i}=a_{i}^{D-d_{i}},$ (1)
accounting for dimension reduction for a subdomain $\Omega_{i}$ with any dimension $d_{i}\leq D$.
@@
Figure 1: Illustration of the mixed-dimensional geometry. The higher-dimensional subdomain $\Omega_{h}$ is coupled to the lower-dimensional subdomain $\Omega_{l}$ through the interfaces $\Gamma_{j}$ and $\Gamma_{k}$. Note that $\partial_{j}\Omega_{h},\partial_{k}\Omega_{h},\Gamma_{j},\Gamma_{k}$ and $\Omega_{l}$ all coincide geometrically. Figure adapted from [Stefansson_2021]
@@ 2.2 Conservation laws and constitutive relations
Next, we define constitutive relations for the various quantities defined above, most of which are derived from [coussy]. The volumetric fluid flux is modeled by Darcy’s law:
@@
while in fracture subdomains, it is given by a cubic law [permeability]:
$\bm{\mathcal{K}}_{i}=\frac{\mathcal{A}_{i}^{2}}{12}\mathbf{I}_{D-1}\ \ ,\ \ d_{i}=D-1,$ (6)
where $\mathcal{A}_{i}$ denotes the hydraulic aperture, which we distinguish from the volumetric aperture $a_{i}$. Constitutive laws for both apertures will be defined in section 2.3, after some new notation has been introduced. Finally, intersection permeability is computed as the average permeability of the intersecting fractures. Solid density is assumed to be constant, while the density of the slightly compressible fluid is given by (we omit the subscript $i$ for readability):
@@
where $\gamma$ is the compressibility of the fluid. The subscript “ref” is used to denote reference values of variables. The interface flux $v_{j}$ is given by an averaged Darcy’s law [martin]:
@@
Here, $\bm{n}_{h}$ is the outward normal vector of the higher-dimensional neighboring subdomain. The interface permeability $\mathcal{K}_{j}$ is also modeled by a cubic law, inheriting the hydraulic aperture from its lower-dimensional neighboring subdomain,
$\mathcal{K}_{j}=\frac{\Pi_{j}^{l}\mathcal{A}_{l}^{2}}{12},$ (9)
@@
where $\rho^{s}$ is the solid density.
@@ 2.3 Fracture contact mechanics
The equations of contact mechanics are imposed on pairs of $D$-dimensional matrix subdomains $\Omega_{h}$ and D1D-1-dimensional fracture subdomains $\Omega_{l}$ that are connected through two interfaces $\Gamma_{j},\Gamma_{k}$, as illustrated in figure 1. Some new variables are introduced at this point, namely the fracture contact traction $\bm{\lambda}_{l}$, defined on the fracture domain $\Omega_{l}$, and the displacements $\bm{u}_{j},\bm{u}_{k}$ of the interfaces $\Gamma_{j},\Gamma_{k}$; these are treated as additional primary unknowns to be solved for. For readability purposes, we will omit the subscript $l$ from the contact traction. The forces on the two fracture surfaces must be balanced according to Newton’s third law. Taking into account that the pressure from the fluid in the fracture contributes to the total fracture traction, the following balance of forces are imposed at the interfaces:
@@
$\begin{cases}\text{If}\ \lambda_{n}=0\implies\bm{\lambda}_{\tau}=0.\\ \text{Else}:\\ \ \ \begin{aligned} \lVert\bm{\lambda}_{\tau}\rVert&\leq-\mu\lambda_{n},\\ \lVert\bm{\lambda}_{\tau}\rVert&<-\mu\lambda_{n}\implies[[\dot{\bm{u}}]]_{\tau}=\bm{0},\\ \lVert\bm{\lambda}_{\tau}\rVert&=-\mu\lambda_{n}\implies\exists\ \zeta\in\mathbb{R}^{+}:[[\dot{\bm{u}}]]_{\tau}=\zeta\bm{\lambda}_{\tau},\end{aligned}\end{cases}$ (18)
where $\mu$ is the coefficient of friction and $\dot{\bm{u}}]]_{\tau}$ is the time derivative of the tangential displacement jump. We will for convenience refer to the entirety of equation (18) as the Coulomb friction law. Finally, we define the constitutive laws for the volumetric and hydraulic apertures. In fracture subdomains, these apertures are set to equal the normal component of the interface displacement jump, with additional constant reference apertures $a_{\mathrm{ref}},\mathcal{A}_{\mathrm{ref}}$ taking into account small-scale roughnesses in the undeformed state:
$a_{i}=a_{\mathrm{ref}}+[[\bm{u}]]_{n}\ \ ,\ \ d_{i}=D-1,$ (19)
$\mathcal{A}_{i}=\mathcal{A}_{\mathrm{ref}}+[[\bm{u}]]_{n}\ ,\ \ d_{i}=D-1.$ (20)
For intersection subdomains, the apertures are computed as the average apertures of the intersecting fractures:
$a_{i}=\frac{1}{\lvert\hat{S}_{i}\rvert}\sum_{j\in\hat{S}_{i}}\Xi_{j}^{i}\Pi_{j}^{h}a_{h}\ \ ,\ \ d_{i}<D-1.$ (21)
$\mathcal{A}_{i}=\frac{1}{\lvert\hat{S}_{i}\rvert}\sum_{j\in\hat{S}_{i}}\Xi_{j}^{i}\Pi_{j}^{h}\mathcal{A}_{h}\ \ ,\ \ d_{i}<D-1.$ (22)
The volumetric aperture $a_{i}$ enters the fracture flow model through the specific volume (1), while the hydraulic aperture $\mathcal{A}_{i}$ governs the fracture permeability through the cubic law (6). Both cases result in a strong nonlinear coupling. We hypothesize a strong impact of the degree of coupling and nonlinearity on the performance of the nonlinear solvers. This will be investigated in the numerical experiments of section 4, by a systematic comparison of different linearizations of the apertures, where either one or both of the volumetric and hydraulic apertures are kept fixed to reduce the degree of nonlinearity.